# Data-Driven Loss Functions for Inference-Time Optimization in Text-to-Image Generation

### v1

@@ 5.2 Qualitative Analysis
Figures 4 and 6 show qualitative examples across four base models and several baselines. Compared to other methods, our approach consistently generates high-quality images that accurately reflect the spatial relations in the prompts. It also handles generating atypical and uncommon scenarios well, as can be seen in  Fig. 4 with “A bus below a toothbrush” and in  Fig. 6 with “An elephant below a surfboard. In contrast, other baselines tend to neglect entities: in Fig. 4, none depict the vase from “A vase above a fire hydrant”, and the zebra or the chair are missing from “A chair to the left of a zebra”. These examples, along with Figure 5, demonstrate how our method overcomes the three main failure modes: incorrect object placement, entity neglect, and object fusion.
@@
In Appendix B, qualitative comparisons with COMPASS show we achieve equivalent spatial accuracy while maintaining style and aesthetics closer to the base model. For additional qualitative results, see Appendix  C, which includes extensive comparisons on both GenEval (Figures 11, 12, 13, and 14) and T2I-CompBench (Figures 15, 16, 17, and 18) benchmarks, as well as uncurated generation results in Fig. 10.
@@ 5.3 Ablation Study
Leakage Handling w/o dual inversion (pos only) 0.425
@@
We examine the key design choices of our method through systematic ablation on FLUX.1-schnell. Results are summarized in Table 3.
Leakage Handling. Our dual-inversion approach is crucial for addressing relation leakage. Without it, training on positive prompts only yields 0.42.5. The dual-inversion strategy eliminates this leakage, enabling our method to reach 0.57.
@@ 6 Limitations
Figure 7: Method Limitations. (a) 2D spatial relations may not align with 3D scene geometry—the spoon appears below the bear in 2D but is in front of it in 3D. (b) Multiple spatial relationships with three relations: When prompted with four objects and three spatial relations, our method is sometimes able to generate matching images. However, typical failure cases—such as entity neglect—still occur
Our method has several limitations. First, we train on 2D spatial relations, which can conflict with 3D scene understanding. As shown in Fig. 7(a), a spoon positioned in front of a bear in 3D space may appear ”below” it in the 2D projection.
Second, handling scenarios with multiple spatial relationships becomes more challenging when three relations are involved compared to two. Our accuracy drops from 0.28 with two relations to just 0.12 with three, showing that as the number of relations increases, accuracy tends to decrease.  Fig. 7(b) shows both successful generations and typical failure cases, such as entity neglect.
Finally, despite substantial improvements, there is still a lot of headroom for improvement. Entity neglect and incorrect positioning persist, as shown in uncurated generation results in Fig. 10 in the Appendix.
@@ 7 Conclusions
We presented Learn-to-Steer, a data-driven alternative to handcrafted spatial losses for test-time steering of text-to-image diffusion models. By training a lightweight classifier to decode spatial relationships directly from cross-attention maps and deploying it as a learned loss function, we achieve substantial improvements in spatial accuracy across multiple models without any fine-tuning. A key advantage of our approach is that it preserves the model’s full capabilities by applying steering only when spatial relations are needed, avoiding the performance degradation often seen with fine-tuning methods. While test-time optimization does increase inference time, this selective application ensures that the model’s general generative abilities remain intact for prompts without spatial constraints. This cost enables applications requiring high spatial accuracy—including generating multiple relations data. Our work demonstrates that learning from the model’s own internal representations can be more effective than handcrafted objectives, opening new directions for improving text-to-image alignment.
@@ Acknowledgments
A Additional Implementation Details
A.1 Adapting FOR’s handcrafted spatial loss to Flux-based models
@@
A.2 Benchmarks
To ensure a fair comparison, we normalize all relation names to {”above”, ”below”, ”to the left of”, ”to the right of”} across benchmarks. All baselines we reproduced using this normalized set. We will release normalized test sets for both benchmarks.
A.3 Multiple Relations Evaluation
@@
We investigated why current benchmarks do not support automatic verification of multiple relations. Specifically, we focused on the GenEval [14] framework and identified the main culprit as the object detector, which we replaced with CLIPSeg [34]. To enable evaluation, we generated metadata files in GenEval’s expected structure, where each prompt is paired with a list of required objects along with their pairwise relations. To verify that our automatic evaluation is reliable, we compared its results against manual evaluation conducted by one of the authors (Table 4). The scores align closely across all settings.
A.3.1 Evaluation Objects
For the multi-relation evaluation, we used the following 17 object categories (inspired by Textual-Inversion [12] and DreamBooth [41]): backpack, bowl, cat, sneakers, duck, robot, teapot, vase, corgi, mug, furby, teddy bear, frog, mouse, table, chair, and umbrella.
A.4 Architecture Details
@@
A.5 Other details
@@
B Comparison with COMPASS
While fine-tuning methods like COMPASS [50] achieve strong performance on spatial relations, they often come at the cost of degrading other alignment capabilities. Table 5 shows this trade-off using results reported in the COMPASS paper.
Model Position ↑ Count ↑ Color ↑
Vanilla SD2.1 0.07 0.44 0.85
COMPASS 0.51 0.20 0.71
Learn-to-Steer (Ours) 0.54 0.44* 0.85*
As shown in Table 5, COMPASS’s spatial relation improvements come with measurable degradation in other capabilities: color accuracy drops from 0.85 to 0.71 and counting accuracy drops from 0.44 to 0.20. In contrast, our test-time optimization approach preserves the original model’s capabilities across all metrics. Additionally, the qualitative comparison in Fig. 8 shows that our method achieves equivalent spatial accuracy while maintaining style and aesthetics closer to the base model.
C Additional Qualitative Results
We perform extensive comparisons on both GenEval (Figures 11, 12, 13, and 14) and T2I-CompBench (Figures 15, 16, 17, and 18).

### v2

@@ 5.2 Qualitative Analysis
Figure 4 (also Suppl Fig. 14) show qualitative examples across four base models and several baselines. Compared to other methods, our approach consistently generates high-quality images that accurately reflect the spatial relations in the prompts. It also handles generating atypical and uncommon scenarios well, as can be seen in Fig. 4 with “A bus below a toothbrush” and in Fig. 14 with “An elephant below a surfboard. In contrast, other baselines tend to neglect entities: in Fig. 4, none depict the vase from “A vase above a fire hydrant”, and the zebra or the chair are missing from “A chair to the left of a zebra”. These examples, along with Fig. 13 (Appendix), demonstrate how our method overcomes the three main failure modes: incorrect object placement, entity neglect, and object fusion.
Interpretability:  Fig. 5 visualizes cross-attention maps throughout denoising, showing how the model establishes spatial relations of two entities. Initially ($t=T$), attention maps are diffuse and overlapping. During denoising, they gradually separate - corgi localizing left, teapot right - matching the prompt. By $t=0$, objects are clearly disentangled with correct spatial alignment.
Figure 5: Evolution of cross-attention maps across denoising steps. Maps shown are after the steering step (See Fig. 3). At early steps ($t=T$), maps are diffuse and overlapping. During denoising, the corgi localizes on the left and the teapot on the right, converging to the intended spatial arrangement by $t=0$.
5.3 Generalization and Scalability
We evaluate our method’s generalization capabilities across multiple dimensions: out-of-distribution categories, compositional complexity, dense prompts, personalized models, and diagonal spatial relations.
Generalization to unseen categories: We test Learn-to-Steer on object categories never seen during training. This out-of-distribution (OOD) evaluation uses 200 prompts across 16 object categories unseen by both our method and COMPASS [55]: Only 6 categories from T2I-CompBench were external to both methods, and 10 additional categories were proposed by ChatGPT, ensuring no training overlap. See Sec. C.5 for a full list. Table 2 shows that Learn-to-Steer maintains strong performance on unseen object categories, achieving 0.60 on FLUX.1-dev and outperforming COMPASS on SD models (0.60 vs. 0.52 on SD2.1).
Compositional scaling to multiple relations: We evaluate how Learn-to-Steer generalizes to prompts with multiple spatial relations – a scenario never seen during training. We created 200 multi-relation prompts, ranging from two relations among three objects to complex cases with 4–5 objects and up to three relations, using both in-distribution (from the training set) and OOD categories as described above. We extended GenEval to support automatic verification of multiple relations and added open-vocabulary evaluation to recognize our new categories (details in Sec. C.4). Results are shown in Table 3. FLUX.1-dev performs poorly (accuracy $\leq$ 0.05); COMPASS, a training-based method, degrades severely with multiple relations (accuracy $\leq$ 0.07 with 3 relations), despite strong single-relation performance. This sharp decline reflects the limitations of training-based approaches outside their training distribution. In contrast, without any new training, Learn-to-Steer achieves substantial gains: for in-distribution categories with 2 relations, its accuracy of 0.36 on 3-4 objects matches the expected independent probability $(0.61)^{2}=0.372$. Accuracy remains strong for OOD categories. These results highlight that our test-time optimization scales better with the number of relations, as performance degrades gracefully, even for OOD categories. See Fig. 16 (Appendix) for qualitative comparisons.
Beyond four basic spatial relations: To evaluate generality, we further trained a classifier for diagonal spatial relations (top-left, bottom-left, top-right, bottom-right). Since GenEval lacks an automated verifier for these relations, we could not use it to generate synthetic data for training the classifier, and therefore trained using only GQA data (see Sec. C.6). The classifier achieved similar performance (38%) to the setting of non-diagonal relations that were trained on GQA alone (Tab. 4, “Real images only (GQA)”). See qualitative examples in Fig. 11 (Appendix).
Zero-shot transfer to personalized models: Learn-to-Steer works out-of-the-box with personalized LoRA models - establishing the spatial relations jointly with the personalized content.  Fig. 6, 12 show two subjects - Eric the cat [16], Margot Robbie [42], and a style - Retro Pixel [41].
Figure 6: Zero-shot transfer to personalized subjects and style LoRAs from HuggingFace. More results in Fig. 12 (Appendix)
Dense Prompts: Fig. 9 (Appendix) shows that our approach maintains spatial correctness under dense prompts enhanced with attributes and scene details.
@@ 5.3 Generalization and Scalability
In Appendix A.1, qualitative comparisons with COMPASS show we achieve equivalent spatial accuracy while maintaining style and aesthetics closer to the base model. Additional results in Appendix D include extensive comparisons on both GenEval (Figures 17, 18, 19, and 20) and T2I-CompBench (Figures 21, 22, 23, and 24).
@@ 5.4 Ablation Study
Leakage Handling w/o leakage handling 0.425
@@
We examine the key design choices of our method through systematic ablation on FLUX.1-schnell. Results are summarized in Table 4. See Appendix B for additional ablations.
Leakage Handling. Our augmentation strategy, using mismatched prompts for inversion is crucial for addressing relation leakage. Without it, training on positive prompts only yields 0.425. The mismatched-prompt strategy eliminates this leakage, enabling our method to reach 0.57.
@@ 6 Limitations
Figure 7: Method Limitations. (a) 2D spatial relations may not align with 3D scene geometry—the spoon appears below the bear in 2D but is in front of it in 3D. (b) Multiple spatial relationships with three relations: When prompted with four objects and three spatial relations, our method can generate matching images. However, typical failure cases, like entity neglect still occur.
Our method has several limitations. First, we train on 2D spatial relations, which can conflict with 3D scene understanding. As shown in Fig. 7(a), a spoon positioned in front of a bear in 3D space may appear ”below” it in the 2D projection. Second, handling scenarios with multiple spatial relationships becomes more challenging as the number of relations grow.  Fig. 7(b) shows both successful generations and typical failure cases, such as entity neglect.
Finally, despite substantial improvements, there is still a lot of headroom for improvement. Entity neglect and incorrect positioning persist, as shown in the uncurated generation results in Fig. 15 in the Appendix.
@@ 7 Conclusions
We presented Learn-to-Steer, a data-driven alternative to handcrafted spatial losses for test-time steering of text-to-image diffusion models. By training a lightweight classifier to decode spatial relationships from cross-attention maps and using it as a learned loss, we achieve substantial spatial accuracy improvements across multiple models without fine-tuning. A key advantage of our approach is that it preserves the model’s full capabilities by applying steering only when needed, avoiding the performance degradation often seen with fine-tuning methods. While test-time optimization increases inference time, this selective application ensures that the model’s general generative abilities remain intact for prompts without spatial constraints. This cost enables applications requiring high spatial accuracy - including generation of multiple relations data. Our work demonstrates that learning from the model’s internal representations surpasses handcrafted objectives, opening new directions for improving text-to-image alignment.
@@ Acknowledgments
A Comparison with Additional Methods
A.1 Comparison with COMPASS
While our method demonstrates strong performance across various models and benchmarks, fine-tuning methods like COMPASS [55] can achieve higher accuracy in specific in-distribution settings, though often at the cost of degrading other alignment capabilities. For instance, on the GenEval benchmark with SD 1.4, COMPASS achieves a spatial accuracy of 43–46% compared to our 36% (Table 1). We attribute this to the nature of fine-tuning, which modifies model weights and can lead to specialization on the training distribution. A more fine-grained analysis suggests this specialization may be narrow. On T2I-CompBench, which uses stricter geometric criteria, our method outperforms COMPASS for all models (e.g., 0.278 vs. 0.254 on SD 1.4); additionally, Table 6 shows that COMPASS’s advantage on SD 1.4 is largely driven by a significant performance gap on the “left of” relation (59.2% vs. 40.8%), while performance on other relations is more comparable.
Importantly, COMPASS’s specialization appears to come at a cost to generalization, particularly in complex, out-of-distribution scenarios (Tables 2, 3). This is most evident when handling prompts with multiple spatial relations—a compositional challenge not seen during training. As shown in Table 3, our method’s performance degrades gracefully, whereas COMPASS’s accuracy drops sharply. For instance, on prompts with three objects and two relations involving OOD categories, our method achieves 28% accuracy compared to COMPASS’s 12%. The performance gap widens with complexity: for four objects and three relations with OOD categories, our method maintains 19% accuracy, while COMPASS drops to just 5%. This highlights the advantage of our test-time optimization in maintaining compositional structure and degrading gracefully under complexity.
This trade-off extends to other capabilities beyond spatial reasoning. Our method better preserves the original aesthetics and style of the base model, as demonstrated in the qualitative comparison in Fig. 8. We quantify this observation in Table 5, which shows that images generated with our method have a higher visual similarity to the base model’s generations using both DINO [39] and CLIP [43] similarities (e.g., on FLUX.1-dev: 0.863 vs. 0.775 for DINO). Furthermore, as shown in Table 7, COMPASS’s improvements in spatial control are accompanied by degradation in other areas: on SD2.1, color accuracy drops from 0.85 to 0.71, and counting accuracy falls from 0.44 to 0.20. In contrast, our test-time approach is applied selectively only when spatial relations are present in the prompt, thereby preserving the base model’s performance on these dimensions.
Base Model Methods DINO ↑ CLIP ↑
Flux-dev
Base vs. COMPASS 0.7751 0.7721
Base vs. Ours 0.8625 0.8767
SD 1.4
Base vs. COMPASS 0.7135 0.7530
Base vs. Ours 0.7360 0.8012
Table 5: Our approach generates images that are closer to the base model. In Figure 8 we see that qualitatively style and aesthetics are closer to the base model.
FLUX.1-dev SD 1.4
Relation Ours COMPASS Ours COMPASS
above 61.5% 63.5% 30.8% 39.4%
below 67.0% 54.5% 33.9% 38.4%
left of 60.5% 56.6% 40.8% 59.2%
right of 54.6% 67.6% 40.7% 40.7%
Table 6: Per-relation accuracy comparison. While FLUX.1-dev results are comparable between methods, in SD 1.4 COMPASS shows notably higher accuracy for “left of” (59%) compared to other relations.
Figure 9: Qualitative results with dense prompts. Our method preserves the intended spatial relations even when objects are described with attributes and embedded in detailed scene contexts.
Model Position ↑ Count ↑ Color ↑
Vanilla SD2.1 0.07 0.44 0.85
COMPASS 0.51 0.20 0.71
Learn-to-Steer (Ours) 0.54 0.44* 0.85*
A.2 Comparison with Layout-guided Methods
Table 8 compares Learn-to-Steer with layout-guided approaches, such as RAGD [10], 3DIS [11], and MIGC [58]. Results for these methods are sourced from RAGD [10], where GenEval’s performance was evaluated using MLLM-generated layouts. Comparisons against MLLM-driven pipelines are inequitable due to their massive parameter counts and undisclosed training data.
Method G. Pos. ↑
3DIS† [11] 0.72
MIGC† [58] 0.76
RAGD† [10] 0.80
Learn-to-Steer (Ours) 0.61
Table 8: Comparison of spatial relation scores from the GenEval benchmark (G. Pos.) against layout-guided methods. †  results are reported from [10]. Note that unlike Learn-to-Steer, these methods rely on explicit layout inputs.
B Additional Ablation Studies
Aggregation Variant G. Pos. ↑
Full Concatenation (Ours) 0.7
Mean: Per-Layer 0.69
Mean: Global (All layers/heads) 0.78
Table 9: Ablation results for different attention aggregation strategies on FLUX.1-dev (GenEval validation set).
Figure 10: Impact of Aggregation on Visual Aesthetics. Comparing our chosen method (Full Concatenation) against the Mean Aggregation baselines. While Mean Aggregation forces the objects into position, it frequently results in loss of background detail and vividness. Our method preserves the visual richness of the base model.
The inputs to the relation classifier are the cross-attention maps of two objects, alongside the current denoising timestep. We analyzed how different aggregation strategies for the attention maps affect performance. Specifically, we compared three approaches: (1) Mean (Global): Computing the mean across all layers and heads; (2) Mean (Per-Layer): Computing the mean across heads within each layer, then concatenating the layer outputs; and (3) Full Concatenation: Retaining all maps by concatenating all layers and heads into a single feature vector.
While Table 9 demonstrates that the “Mean (Global)” strategy achieves the highest alignment score (0.78), we find that it comes at a cost to image quality, consistently yielding over-constrained results with “dull” or neutral backgrounds (See Fig. 10). Similarly, the “Mean (Per-Layer)” strategy, which scores slightly lower than our method (0.69 vs. 0.70), also exhibits inconsistent visual quality, occasionally suffering from the same background degradation. In contrast, the “Full Concatenation” strategy provides the most balanced setup, offering strong adherence to the spatial prompt while maintaining high visual aesthetics. Therefore, we decided to use this strategy for our final method.
C Additional Implementation Details
C.1 Adapting FOR’s handcrafted spatial loss to Flux-based models
@@
C.2 Benchmarks
To ensure a fair comparison, we normalize all relation names to {”above”, ”below”, ”to the left of”, ”to the right of”} across benchmarks. All baselines we reproduced use this normalized set. We will release normalized test sets for both benchmarks.
C.3 OOD Evaluation Details
Our out-of-distribution (OOD) evaluation tests generalization on 200 prompts across 16 object categories unseen by both our method and COMPASS [55]. We identified 6 categories from T2I-CompBench vocabulary that were out-of-distribution for both methods, and used ChatGPT to generate 10 additional diverse categories, ensuring no training data overlap.
C.4 Extending GenEval to an Open-Set of Categories and Multiple Relations
The current GenEval benchmark is constrained for only recognizing the 80 COCO categories, and can only verify a single pair of relations. To facilitate an open set of object categories, we replaced the object detector with CLIPSeg [35], a zero-shot object segmentation approach. To evaluate multiple relations, we generated metadata files in GenEval’s expected structure, where each prompt is paired with a list of required objects along with their pairwise relations, verify each relation, and indicate success only when all relations are correctly generated in an image. To verify that our automatic evaluation is reliable, we compared its results against manual evaluation conducted by one of the authors on a mixed set of in-distribution and OOD categories (Table 10). The scores align closely across all settings.
@@ C.4 Extending GenEval to an Open-Set of Categories and Multiple Relations
C.5 Object Categories for Multi-Relation Evaluation
For the multi-relation evaluation, similar to the single-relation case, we tested on both in-distribution and out-of-distribution (OOD) object categories. The specific objects used were:
In-distribution categories (16 objects): backpack, bowl, cat, skateboard, bottle, cake, tie, vase, sheep, toothbrush, teddy bear, suitcase, mouse, knife, chair, and umbrella.
Out-of-distribution categories (16 objects): teapot, corgi, furby, robot, turtle, rabbit, butterfly, key, pig, bee, tiger, hammer, monkey, razor, snake, and stapler.
C.6 Extending Our Method to Diagonal Spatial Relations
Since there is no labeled data for diagonal relations (top-left, top-right, bottom-left, bottom-right) readily available, we generated training samples by inferring relations directly from GQA bounding boxes. For each image, we examined all object pairs and computed their relative offsets and angles. Pairs whose center displacement fell within diagonal sectors were labeled accordingly, while cases with large overlaps were discarded.
Using these labels, we constructed cross-attention training samples following the same procedure as for the four basic relations (above, below, left, right). For a given relation (e.g., “top-left”), we inverted the image twice: once with the correct relation and once with an incorrect relation sampled uniformly from the remaining options (e.g., “bottom-left”). This shows that our dual-image inversion strategy can naturally extend to more complex spatial relations such as diagonals.
We then trained a diagonal relation classifier in the same way as for the four basic relations, with one exception: as a quick prototype of the aggregation module, we computed the average across all cross-attention maps before forwarding the representation to the classification module.
C.7 Potential limitations of dual-inversion
Our dual inversion strategy has an underlying assumption that the inversion process succeeds, and generates well-behaving attention maps even for the incorrect prompt. While we did not experiment with inverting very complex prompts, inversion it is likely to be less stable in these cases. Another limitation arises when inverting images that contain multiple instances of the same object (2 cats and a dog). Such images and prompts are known to be less stable.
Figure 11: Qualitative comparison of diagonal spatial relations. The base model often misplaces objects or ignores the diagonal specification, while our method better adheres to the intended spatial layout.
C.8 Dense Prompt Construction
To create dense prompts, we asked ChatGPT [38] to enhance base prompts from the GenEval dataset, and generate an enriched variant. Each variant preserved the original spatial relation, added an attribute to every object, placed the objects in a realistic location (e.g., kitchen, living room, park), and included a vivid description of the location with detailed context.
C.9 Architecture Details
@@
C.10 Other details
Relation Extraction. Relation triplets are extracted from the prompt using SpaCy [21] a natural lang. parsing library, with some hard-coded exceptions, like ”fire hydrant”.
@@ C.10 Other details
D Additional Qualitative Results
Figure 14 provides comprehensive comparisons across all four base models on GenEval prompts, demonstrating consistent improvements in spatial accuracy. Figure 13 (Appendix), demonstrate how our method overcomes the three main failure modes: incorrect object placement, entity neglect, and object fusion.
We perform extensive comparisons on both GenEval (Figures 17, 18, 19, and 20) and T2I-CompBench (Figures 21, 22, 23, and 24).
Figure 12: More qualitative results of zero-shot transfer to personalized LoRAs of both subject and style.